What is the equation of a circle with center (5,-7) and diameter 6?
Found 2 solutions by stanbon, nerdybill:
Answer by stanbon(75887)

What is the equation of a circle with center (5,-7) and diameter 6?
-----
radius = 3
----
Equation:
Let (x,y) be any point on the circle.
(x-5)^2 + (y+7)^2 = 3^2
=============================
Cheers,
Stan H.

Answer by nerdybill(7384)

What is the equation of a circle with center (5,-7) and diameter 6?
.
Standard form of a circle is:
(x-h)^2 + (y-k)^2 = r^2
where
(h,k) is the vertex or center
and
r is the radius
.
if diameter is 6
radius (r) is 6/2 or 3
.
Plug into circle equation:
(x-5)^2 + (y-(-7))^2 = 3^2
(x-5)^2 + (y+7)^2 = 9